6. One form of nuclear radiation, beta decay, occurs when a neutron changes into a proton, an electron, and a neutral particl thank u so much!
Answer #1

ł the n 34 - ptt Het ē t de 21 He ap 2n in a) charge in beta le conservation - decay process of trituim nucles 3 Ho 3 Het è tsight hand side: Co (We) Total charge = (2+) + (2-) + (6) = 4+ left hand side total charge = (1+) + (20) = 1 + (P) ) This totLLIURETIRO CULTURE in onsor: Radius of nudeas. this op case, mo rx mo for minimum speed - - » - 6.6481094 200 m 4X7 X 1.5 110